# Fourfold path to Thermality: Inequivalent purifications of Rindler wedge

### v1

@@ V.1 Two-Dimensional CFT on Rindler Spacetime in the Minkowski Vacuum
From Eqs. (3), (4),  (66) and  (188) and after a simple calcualtion the expectation value of the energy-momentum tensor in Rindler $\mathrm{R_{1}}$ gives,
@@
$\begin{split}\langle 0_{M}|:T_{u_{3}u_{3}}:|0_{M}\rangle=\bigg(\frac{\partial_{u_{1}}}{\partial_{u_{3}}}\bigg)^{2}\langle 0_{M}|:T_{u_{1}u_{1}}:|0_{M}\rangle\\ -\frac{\hbar}{24\pi}S(u_{1},u_{3}),\end{split}$ (193)
And,
$\begin{split}\langle 0_{M}|:T_{v_{3}v_{3}}:|0_{M}\rangle=\bigg(\frac{\partial_{v_{1}}}{\partial_{v_{3}}}\bigg)^{2}\langle 0_{M}|:T_{v_{1}v_{1}}:|0_{M}\rangle\\ -\frac{\hbar}{24\pi}S(v_{1},v_{3}),\end{split}$ (194)
Eqs. (193) and  (194) are known as the Virasoro anomaly.
By using Eqs. (9), (66),Eqs. (189) and Eqs. (193), we get the expectation value in $\mathrm{R_{3}}$ through the excited state $\mathrm{R_{1}}$ is,
$\langle 0_{M}|:T_{u_{3}u_{3}}:|0_{M}\rangle=\frac{\hbar a^{2}}{48\pi},$ (195)
And similarly, using Eqs. (10), (66),Eqs. (190) and Eqs. (194), we get the expectation value in $\mathrm{R_{3}}$ through the excited state $\mathrm{R_{1}}$ is,
$\langle 0_{M}|:T_{v_{3}v_{3}}:|0_{M}\rangle=\frac{\hbar a^{2}}{48\pi},$ (196)
Now, as the above calculation, the Null shifted wedges $\mathrm{R_{2}}$ and $\mathrm{R_{4}}$ through $\mathrm{R_{1}}$ as shown in Figs. 1 and  2 simply gives,
$\langle 0_{M}|:T_{u_{2}u_{2}}:|0_{M}\rangle=\frac{\hbar a^{2}}{48\pi}=\langle 0_{M}|:T_{v_{3}v_{3}}:|0_{M}\rangle,$ (197)
And
$\langle 0_{M}|:T_{u_{4}u_{4}}:|0_{M}\rangle=\frac{\hbar a^{2}}{48\pi}=\langle 0_{M}|:T_{v_{4}v_{4}}:|0_{M}\rangle,$ (198)
[…]
$\langle 0_{M}|:T_{u_{3}u_{3}}:|0_{M}\rangle=\frac{\hbar a^{2}}{48\pi}=\langle 0_{M}|:T_{v_{3}v_{3}}:|0_{M}\rangle,$ (199)
Our analysis shows that the Minkowski vacuum, when viewed from any of the Rindler wedges, whether spatially separated or null shifted, exhibits identical expectation values for the chiral components of the energy–momentum tensor. In every case, the Schwarzian derivative associated with the coordinate transformation reproduces the standard thermal stress tensor, reflecting the Unruh temperature perceived by uniformly accelerated observers. We conclude that regardless of whether the transformation between wedges, demonstrating that all these wedges are conformally related to Minkowski spacetime.
V.2 Two-Dimensional CFT on Rindler Spacetime in the Rindler Vacuum
In this section, we again investigate the expectation values of the energy–momentum tensor in the vacuum state as seen from the Rindler wedge $\mathrm{R_{1}}$, evaluated via four paths.
An important conceptual difference arises immediately, unlike the Minkowski vacuum, the Rindler vacuum $|0_{R_{1}}\rangle$ is not a globally defined Hadamard state on Minkowski spacetime[22]. It is only well-defined within the wedge $\mathrm{R_{1}}$ and not defined on the Horizon, because it is singular on the Horizon, and it is not invariant under translations, boosts, or null shifts. Consequently, extending $|0_{R_{1}}\rangle$ into the null-shifted wedges $\mathrm{R_{2}}$, $\mathrm{R_{4}}$ or $\mathrm{R_{3}}$ is not a symmetry operation of quantum field theory and cannot generically define a consistent state in those regions.This subtlety is crucial for understanding the path dependence that emerges below.
The expectation value of the energy–momentum tensor along Path-1, which maps Minkowski space directly to the Rindler wedge $\mathrm{R_{3}}$ while bypassing $\mathrm{R_{1}}$,has already been derived in the preceding section,as given in Eqs. (191) and  (192).
Next, we compute the energy–momentum tensor according to Path-2. In this case, the field is evaluated in $\mathrm{R_{3}}$,but the vacuum state is taken to be that of the Rindler wedge $\mathrm{R_{1}}$ and is denoted as $|0_{R_{1}}\rangle$.
@@ V.2 Two-Dimensional CFT on Rindler Spacetime in the Rindler Vacuum
$\langle 0_{R_{1}}|:T_{u_{1}u_{1}}:|0_{R_{1}}\rangle=\langle 0_{R_{1}}|:T_{v_{3}v_{3}}:|0_{R_{1}}\rangle=0,$ (202)
@@
$\langle 0_{R_{1}}|:T_{v_{3}v_{3}}:|0_{R_{1}}\rangle=\frac{\hbar a^{2}}{24\pi}e^{-av_{3}}$
@@
Now, we are evaluating the expectation values of energy-momentum tensor in the null-shifted wedge $\mathrm{R_{2}}$ as shown in Fig. 1 with respect to vacuum state in $\mathrm{R_{1}}$. From Eqs. (17) and  (66), we get
@@
A simple calculation, gives
@@
And similarly we can calculate,
@@
In the above discussion we observe that with respect to Rindler vacuum state $\mathrm{R_{1}}$ the component of the $T_{u_{2}u_{2}}$ is zero and $T_{v_{2}v_{2}}$ is non-zero as discussed above.
Now we evaluate the expectation values of the energy–momentum tensor in the Rindler wedge $\mathrm{R_{3}}$ via Path-3, using the Rindler vacuum defined in $\mathrm{R_{1}}$, the expectation value of energy-momentum tensor is
@@
Eq. (210) is exactly match with the Eq. (203) and hence gives the same limiting condition. Similarly, by the simple calculations, we get
@@
Taken together with the results of Sec. V.1, our analysis reveals a clear structural distinction between the two vacuum choices. For the Minkowski vacuum $|0_{M}\rangle$, all four paths connecting the nested Rindler wedges yield the same thermal expectation values in $\mathrm{R_{3}}$, reflecting the conformal equivalence of the wedges and the universality of the Unruh temperature. In contrast, when the Rindler vacuum $|0_{R_{1}}\rangle$ is used, the four paths no longer agree; Path-1 and Path-2 produce distinct stress–tensor components in $\mathrm{R_{3}}$, while Path-3 and Path-4 agree with each other but differ from Paths 1 and 2. This demonstrates that, unlike the Minkowski vacuum, the Rindler vacuum is sensitive to the choice of intermediate wedge, reflecting its non-invariance under the full modular flow connecting the nested Rindler regions.
The above comparison also highlights an essential physical distinction between the two vacuum choices. When the Minkowski vacuum $|0_{M}\rangle$ is used, all paths consistently reproduce the standard thermal stress tensor associated with the Unruh effect. This reflects a genuinely physical thermality; a uniformly accelerated observer perceives the Minkowski vacuum as a thermal state at the Unruh temperature, in complete agreement with quantum field theory in curved spacetime. In contrast, when the Rindler vacuum $|0_{R_{1}}\rangle$ is employed, the resulting stress-tensor components display path dependence and fail to yield a universal thermal spectrum. The apparent thermality that emerges in this case is therefore not physical but merely an effective or coordinate-induced thermality, arising from the restricted domain of the Rindler vacuum rather than from any true thermal bath. Thus, the Minkowski vacuum leads to genuine Unruh thermality, while the Rindler vacuum generates only a formal, non-physical analogue.
VI Conclusion
We have investigated the emergence of thermal behaviour in two-dimensional Rindler spacetime by analysing a hierarchy of spatially shifted and null-shifted wedge constructions in Minkowski space. By combining Bogoliubov transformations with an independent evaluation of the expectation value of the stress-energy tensor, we have shown that thermal occupation numbers, Gibbs thermality, and entanglement across horizons are logically distinct features whose relation depends sensitively on the underlying geometry and choice of vacuum.
Uniformly accelerated observers associated with different Rindler wedges and characterised by distinct accelerations $a$ and $g$ do not, in general, detect particles with respect to each other’s natural vacuum states. This result demonstrates that relative acceleration alone is insufficient to generate particle production or thermality, and that the Unruh effect relies crucially on the presence of a causal horizon and the accompanying entanglement structure of the Minkowski vacuum.
When the global state is the Minkowski vacuum, thermal occupation numbers detected in a single Rindler wedge arise from tracing over degrees of freedom beyond the horizon. The resulting local state is mixed and Gibbsian, with both left- and right-moving sectors exhibiting thermal occupation. In contrast, null-shifted wedge constructions give rise to thermal occupation numbers in the absence of entanglement between complementary wedges. In these cases, the global and local quantum states remain pure, and thermal behaviour emerges solely from Bogoliubov mixing induced by null translations along Rindler horizons. The thermality is therefore selective, affecting only one chiral sector while the complementary sector remains in its vacuum.
The stress-energy tensor provides an independent and physically robust diagnostic of these distinctions. When evaluated in the Minkowski vacuum, its expectation value is universal and insensitive to the particular sequence of wedge constructions, reflecting genuine horizon-induced thermality. By contrast, when evaluated in the Rindler vacuum, the stress-energy tensor exhibits path-dependent and unphysical behaviour, underscoring that apparent thermal occupation numbers alone do not guarantee a physically meaningful thermal state.
Taken together, these results demonstrate the existence of inequivalent purifications of thermal occupation numbers in Rindler spacetime. They establish explicitly that thermal behaviour need not originate from entanglement or mixedness of the quantum state. In this sense, the null-shifted constructions realise a converse perspective on the Unruh effect, in which thermal occupation arises from kinematical mode mixing rather than from tracing over inaccessible degrees of freedom. More broadly, our findings clarify the distinct roles of observer dependence, causal structure, and vacuum choice in the emergence of thermality in relativistic quantum field theory.
@@ Appendix B Particle number density
The left-hand side of Eq. (225) gives the dimensionless number of particles. So, $I_{1}$ and $I_{2}$ must also be dimensionless. To keep the units consistent in Eq. (226), the factor $\frac{1}{2\pi a}$is included inside the integral. The integral $\int_{\frac{\pi\nu}{a}}^{\infty}\frac{1}{2\pi a}\frac{dz}{z}$ exhibits a logarithmic ultraviolet (UV) divergence arising from contributions of high-frequency modes. This type of divergence is characteristic of quantum field theories in curved spacetime and often reflects the accumulation of an infinite density of states near horizons or boundaries. In such contexts, divergences are frequently formalized via expressions like $\delta(0)$—interpreted as an infinite volume or mode density. $\delta(0)$ emerges from continuous mode normalization and is directly linked to spatial volume [23, Sec.2.3]. So the Integral  (226) can be written after simplification;
@@
Which has the same functional form as the Bose–Einstein factor, with $\delta(0)$ interpreted as the formal infinite spatial volume factor.
Appendix C Purity of the State in Region $R_{2}$
In this appendix, we verify that the quantum state defined in region $R_{2}$,
$|\psi_{R_{2}}\rangle=\left(\sum_{n=0}^{\infty}\sqrt{p_{n}}\,|\overleftarrow{n}\rangle\right)\otimes|\overrightarrow{0}\rangle,$ (237)
is a pure state, using the standard formalism of density matrices. Here, the coefficients $p_{n}$ take the Planckian form
$p_{n}=\frac{e^{-n\beta\Omega}}{Z},\qquad Z=\sum_{n=0}^{\infty}e^{-n\beta\Omega},$ (238)
and satisfy $\sum_{n}p_{n}=1$.
While these coefficients resemble thermal Boltzmann factors, in this context, they originate from the Bogoliubov amplitudes describing the mode mixing between the vacua of two Rindler-like regions $R_{1}$ and $R_{2}$ connected by a null translation. This transformation induces Bogoliubov coefficients whose magnitudes follow a thermal distribution. However, the resultant quantum state in the Hilbert space of $R_{2}$ remains a pure vector rather than a mixed state, since the transformation is unitary and no information is lost.
The corresponding density matrix is
$\rho=|\psi_{R_{2}}\rangle\langle\psi_{R_{2}}|=\sum_{n,m}\sqrt{p_{n}p_{m}}\,|\overleftarrow{n}\rangle\langle\overleftarrow{m}|\otimes|\overrightarrow{0}\rangle\langle\overrightarrow{0}|.$ (239)
This operator is a rank-one projector satisfying $\rho^{2}=\rho$. Hence, the state has unit purity,
$\operatorname{Tr}(\rho^{2})=1,$ (240)
and is therefore pure.
To isolate the left-moving sector, we trace over the right-moving vacuum:
$\overleftarrow{\rho}$ $=\operatorname{Tr}_{\overrightarrow{\mathcal{H}}}(\rho)$
$=\sum_{n,m}\sqrt{p_{n}p_{m}}\,|\overleftarrow{n}\rangle\langle\overleftarrow{m}|\,\operatorname{Tr}(|\overrightarrow{0}\rangle\langle\overrightarrow{0}|)$
$=\sum_{n,m}\sqrt{p_{n}p_{m}}\,|\overleftarrow{n}\rangle\langle\overleftarrow{m}|.$ (241)
This is again a rank-one operator and can be written as
$\overleftarrow{\rho}=|\overleftarrow{\psi}\rangle\langle\overleftarrow{\psi}|,\qquad|\overleftarrow{\psi}\rangle=\sum_{n}\sqrt{p_{n}}\,|\overleftarrow{n}\rangle.$ (242)
Therefore,
$\operatorname{Tr}(\overleftarrow{\rho}^{2})=\langle\overleftarrow{\psi}|\overleftarrow{\psi}\rangle=1,$ (243)
confirming that the reduced state $\overleftarrow{\rho}$ is also pure.
Although the coefficients $p_{n}$ exhibit the form of a thermal distribution, they do not represent classical statistical probabilities obtained by tracing out inaccessible degrees of freedom, as is typical for a thermal or Kubo-Martin-Schwinger (KMS) state. Rather, these coefficients arise from the Bogoliubov transformation, which relates mode bases adapted to distinct regions or observers. The squared Bogoliubov coefficients give rise to this thermal-like occupation pattern. It is important to emphasise that the state constructed above is a pure state on the single-wedge algebra, since both chiral sectors appearing in the mode expansion are supported within the same Rindler wedge. No causally complementary algebra is available with respect to which a partial trace could be defined. Consequently, although one chiral sector exhibits a thermal occupation number distribution, the global state remains pure and does not correspond to a thermal density matrix.
Importantly, since no information is lost in this process and the complementary chiral sector remains in its vacuum state, the overall quantum state remains pure. Thus, while the $p_{n}$ serve as thermal weights, they fundamentally encode quantum amplitudes within a coherent pure state, rather than reflecting classical probabilistic mixtures.
$|\psi_{R_{2}}\rangle=|\overleftarrow{\psi}\rangle\otimes|\overrightarrow{0}\rangle,$ (244)
With all apparent “thermality” being a kinematic feature of the mode decomposition, not a sign of mixedness or entanglement.
Appendix D Purity of the State in Region $R_{3}$ via Path 3
In this appendix, we analyse the density matrix corresponding to the quantum state in region $R_{3}$, obtained along Path 3. Along this path, the left-moving modes acquire Bogoliubov thermal structure during the transition $R_{1}\to R_{2}$, and the right-moving modes become thermally excited during the subsequent shift $R_{2}\to R_{3}$. The resulting state in $R_{3}$ takes the form
$|\psi_{R_{3}}\rangle=\left(\sum_{n=0}^{\infty}\sqrt{p_{n}}\,|\overleftarrow{n}\rangle\right)\otimes\left(\sum_{m=0}^{\infty}\sqrt{q_{m}}\,|\overrightarrow{m}\rangle\right),$ (245)
where the coefficients have a Planckian dependence,
$p_{n}=\frac{e^{-n\beta\Omega}}{Z},\qquad q_{m}=\frac{e^{-m\beta\Omega}}{Z},\qquad Z=\sum_{k=0}^{\infty}e^{-k\beta\Omega}.$
Although these factors resemble thermal Boltzmann weights, in this context, they originate from Bogoliubov amplitudes associated with mode mixing under successive null translations. Their apparent thermality is kinematic, not statistical.
The state $|\psi_{R_{3}}\rangle$ is a normalized product vector in the Hilbert space $\mathcal{H}_{R_{3}}$, and is therefore manifestly pure. The corresponding total density matrix is
$\begin{split}\rho_{R_{3}}&=|\psi_{R_{3}}\rangle\langle\psi_{R_{3}}|\\[4.0pt] &=\sum_{n,m,n^{\prime},m^{\prime}=0}^{\infty}\sqrt{p_{n}p_{n^{\prime}}}\,\sqrt{q_{m}q_{m^{\prime}}}\,|\overleftarrow{n}\rangle\langle\overleftarrow{n^{\prime}}|\otimes|\overrightarrow{m}\rangle\langle\overrightarrow{m^{\prime}}|.\end{split}$ (246)
This operator is a rank-one projector, satisfying $\rho_{R_{3}}^{2}=\rho_{R_{3}}$ and $\operatorname{Tr}(\rho_{R_{3}})=1$, which confirms that the global state is pure.
To examine the individual chiral sectors, we trace over one mode. Since the total state is separable, the reduced density matrix of the left-moving mode is
$\overleftarrow{\rho}$ $=\operatorname{Tr}_{\overrightarrow{\mathcal{H}}}(\rho_{R_{3}})$
$=\left(\sum_{n=0}^{\infty}\sqrt{p_{n}}\,|\overleftarrow{n}\rangle\right)\left(\sum_{n^{\prime}=0}^{\infty}\sqrt{p_{n^{\prime}}}\,\langle\overleftarrow{n^{\prime}}|\right)=|\overleftarrow{\phi}\rangle\langle\overleftarrow{\phi}|,$ (247)
where $|\overleftarrow{\phi}\rangle=\sum_{n}\sqrt{p_{n}}\,|\overleftarrow{n}\rangle$. Thus, $\overleftarrow{\rho}$ is a rank-one projector and represents a pure state.
Similarly, tracing over the left-moving sector gives
$\overrightarrow{\rho}=\operatorname{Tr}_{\overleftarrow{\mathcal{H}}}(\rho_{R_{3}})=|\overrightarrow{\phi}\rangle\langle\overrightarrow{\phi}|,\qquad|\overrightarrow{\phi}\rangle=\sum_{m}\sqrt{q_{m}}\,|\overrightarrow{m}\rangle.$ (248)
Both reduced states are pure and separable, confirming that the total state is unentangled.
Although the number-basis coefficients in each sector exhibit thermal-like distributions, the reduced density matrices are rank-one projectors and therefore not thermal in the Gibbs sense [20]. Importantly, this structure is not a thermofield double (TFD); no entanglement is present between the two chiral sectors. Instead, Eq. (245) represents an unentangled superposition of number states with Bogoliubov-induced thermal profiles in each mode sector. The apparent thermality emerges purely from independent Bogoliubov mode mixing along Path 3, not from tracing over complementary horizons.

### v2

@@ V.2 Path 1
From Eqs. (4), (5), (67) and  (158), and after a simple calculation, the expectation value of the energy-momentum tensor in Rindler $\mathrm{R_{1}}$ gives,
@@
V.3 Path 2
In this section, we again investigate the expectation values of the energy–momentum tensor in the vacuum state as seen from the Rindler wedge $\mathrm{R_{1}}$, evaluated via paths 2,3,4.
We compute the energy–momentum tensor according to Path-2. In this case, the field is evaluated in $\mathrm{R_{3}}$, but the vacuum state is taken to be that of the Rindler wedge $\mathrm{R_{1}}$ and is denoted as $|0_{R_{1}}\rangle$.
@@ V.3 Path 2
$\langle 0_{R_{1}}|:T_{u_{1}u_{1}}:|0_{R_{1}}\rangle=\langle 0_{R_{1}}|:T_{v_{1}v_{1}}:|0_{R_{1}}\rangle=0,$ (165)
@@
$\langle 0_{R_{1}}|:T_{v_{3}v_{3}}:|0_{R_{1}}\rangle=0$
@@
The relevant limits for our problem are $u_{3}\to\infty$ and $v_{3}\to\infty$, since we are interested in the particle content near the horizons. We see that it coincides with the answer in path 1. This shows that close to the horizon, there is no difference between the particle content if it is from the Minkowski vacuum or the Rindler vacuum. The distinction is felt when one is far from the horizon. As can be seen in the above equations, the stress tensor approaches zero very far from the horizons, reflecting the fact that the quantum state is the Rindler vacuum.
V.4 Path 3
We now analyze path 3. We evaluate the expectation values of the energy-momentum tensor in the null-shifted wedge $\mathrm{R_{2}}$ as shown in Fig. 1 with respect to the vacuum state in $\mathrm{R_{1}}$. From Eqs. (18) and  (67), we get
@@ V.4 Path 3
A simple calculation gives
@@
This implies there is no flux of left-moving particles, as obtained previously using Bogoliubov coefficients. We now analyze the other sector. And similarly, we can calculate,
@@
In the above discussion, we observe that with respect to the Rindler vacuum state, $T_{v_{2}v_{2}}$ has the same flux as thermal radiation in the near-horizon limit. Far from the horizon, the flux goes to zero. We again observe that observers close to the horizon observe a thermal flux of left-moving particles.
Now we evaluate the expectation values of the energy–momentum tensor in the Rindler wedge $\mathrm{R_{3}}$ via Path-3, using the Rindler vacuum defined in $\mathrm{R_{1}}$, the expectation value of the energy-momentum tensor is
@@
Eq. (173) is exactly matches with the Eq. (166). We note that in the limit of $u_{3}\to\infty$, the flux matches the thermal flux. So close to the horizon, the stress tensor component indicates thermal distribution of right-moving particles. We now evaluate the particle flux. Similarly, by the simple calculations, we get
@@ V.5 Path 4
VI Conclusions and Discussion
In the article, we showed that there are multiple paths that yield a thermal distribution of particles in a given Rindler wedge, when we focus on particle content close to the horizon. For observers close to the horizons, we see that all four paths yield the same thermal density for both left-moving and right-moving particles. The various paths, based on the analysis in the article, are: a) Minkowski spacetime with scalar field in vacuum state, b) Rindler wedge 1 $R_{1}$ with the scalar field in Rindler vacuum, c) Rindler wedge $R_{2}$ with with a thermal flux of left moving particles d) Rindler wedge $R_{4}$ with a thermal flux of right moving particles. Here are a few questions that can be asked based on the above observations. We note that there is an inherent degeneracy when defining a given path. The degeneracy arises from varying the “distances” along the null and spatial shifts. There are various superset wedges (a few spatially shifted, and a few null shifted with respect to the wedge $R_{3}$ that yield the same particle content close to the horizon in $R_{3}$.
Firstly, can one define a notion of entropy associated with a Rindler wedge with the observed thermal density of particles? How to use this degeneracy to consistently define a notion of information entropy is a question we leave for future consideration.
An important consequence of the result could be for Hawking radiation and the evaporation of a black hole. The qualitative arguments are presented below. The black hole evaporation is modeled as follows. The black hole emits a flux of radiation at a temperature inversely proportional to the mass $M$ of the black hole (when we assume the cosmological constant to be zero). The total mass, therefore, decreases to MdMM-dM. The black hole now radiates at a higher temperature because its mass has decreased. The black hole continues to lose mass until it completely evaporates. If we understand this process in discrete steps. The exterior of the black hole with mass $M$ is a subset of the new exterior of a black hole with mass MdMM-dM. The near-horizon geometry of the black hole with mass $M$ resembles a Rindler wedge, complete with features like thermal flux of particles. We name it $R_{2}$. Now, when we go to the next step in Hawking evaporation, we consider a superset of $R_{2}$, which is again a Rindler wedge in the near-horizon geometry. So the main question is “Which superset does it decay into?” Is it another wedge similar to $R_{2}$ with again a thermal flux of particles? The current models of Hawking radiation assume that the superset again has a thermal flux of radiation. The work presented in this article suggests a different possibility. The superset could be a wedge $R_{1}$ in vacuum. If this is the case, Hawking radiation will temporarily cease when the black hole mass reaches MdMM-dM, since the superset is in a vacuum. Of course, it can be argued that a Rindler vacuum in the current setting is similar to the Boulware vacuum and might not be physically realizable. This decay could be unstable and has to be explored rigorously. If this decay mode is allowed, the continuous evaporation of a black hole might be punctuated by a series of pauses and bursts of radiation.
@@ Appendix B Particle number density
The left-hand side of Eq. (188) gives the dimensionless number of particles. So, $I_{1}$ and $I_{2}$ must also be dimensionless. To keep the units consistent in Eq. (189), the factor $\frac{1}{2\pi a}$is included inside the integral. The integral $\int_{\frac{\pi\nu}{a}}^{\infty}\frac{1}{2\pi a}\frac{dz}{z}$ exhibits a logarithmic ultraviolet (UV) divergence arising from contributions of high-frequency modes. This type of divergence is characteristic of quantum field theories in curved spacetime and often reflects the accumulation of an infinite density of states near horizons or boundaries. In such contexts, divergences are frequently formalized via expressions like $\delta(0)$—interpreted as an infinite volume or mode density. $\delta(0)$ emerges from continuous mode normalization and is directly linked to spatial volume [19, Sec.2.3]. So the Integral  (189) can be written after simplification;
@@
Reflecting the familiar Bose–Einstein distribution at the Unruh temperature, with $\delta(0)$ interpreted as the formal infinite spatial volume factor.
Appendix C Bogoliubov Transformation for two observes in Rindler spacetime
Figure 6: Rindler spacetime with observer having acceleration a and g with green and red respectively.
$T=\frac{e^{a\;x_{1}}}{a}\;\sinh\;(a\;t_{1}),$ (200)
$X=\frac{e^{a\;x_{1}}}{a}\;\cosh(a\;t_{1}).$ (201)
which in light-cone coordinates takes the form:
$U=T-X=-\frac{e^{-a\;u_{1}}}{a},$ (202)
$V=T+X=\frac{e^{a\;v_{1}}}{a}.$ (203)
Similarly, for Rindler coordinates with acceleration $g$, we have:
$T=\frac{e^{g\;x_{2}}}{g}\;\sinh\;(g\;t_{2}),$ (204)
$X=\frac{e^{g\;x_{2}}}{g}\;\cosh(g\;t_{2}),$ (205)
and in light-cone form:
$U=T-X=-\frac{e^{-g\;u_{2}}}{g},$ (206)
$V=T+X=\frac{e^{g\;v_{2}}}{g},$ (207)
$u_{1}=\frac{-1}{a}\;ln\bigg(\frac{a}{g}\bigg)+\frac{g}{a}\;u_{2},$ (208)
$v_{1}=\frac{1}{a}\;ln\bigg(\frac{a}{g}\bigg)+\frac{g}{a}\;v_{2},$ (209)
Substituting Eq. (208) into Eq. (210), we obtain:
By the integration technique, we obtained the following result.
Similarly, the $\beta^{*}$-coefficient becomes
Hence, by integrating Eq. (214), we get
$\overrightarrow{\beta^{*}}_{\Omega,\omega}=0,$ (215)
Following a similar analysis for left-moving modes, we obtain:
$\overleftarrow{\beta^{*}}_{\Omega,\omega}=0,$ (217)